# How Breakable Is Privacy: Probing and Resisting Model Inversion Attacks in Collaborative Inference

### v4

@@ 5.3. The Design of Loss Function
Some defense approaches utilize Constrained Covariance (COCO) (Gretton and Herbrich, 2005) to estimate correlation, focusing primarily on the linear relationship between $x$ and $z$. However, since neural networks predominantly perform nonlinear operations, studies have shown that this constraint is not only ineffective but may also lead to reduced usability (Peng et al., 2022). Alternatively, some approaches employ the Hilbert-Schmidt Independency Criterion (HSIC) (Gretton et al., 2005) to estimate nonlinear correlation. However, kernel-based calculations in HSIC rely heavily on the appropriate selection of $\sigma$. Improper $\sigma$ value can introduce bias in correlation estimation, limiting the generalizability of this method and reducing its effectiveness across various scenarios.
@@
(12) $\mathcal{L}_{\text{d Cor}}=\mathrm{d}\operatorname{Cor}(x,z)=\frac{\mathrm{d}\operatorname{Cov}(x,z)}{\sqrt{\mathrm{d}\operatorname{Var}(x)\cdot\mathrm{d}\operatorname{Var}(z)}}$
@@
To compute the numerator, a distance matrix $A$ is first calculated, where each element is the Euclidean distance between samples, defined as $A=\left\|x_{i}-x_{j}\right\|_{2}$. To remove the influence of sample shifts, the distance matrix is centered as $A_{ij}^{c}=A_{ij}-\overline{A_{i}}-\overline{A_{j}}+\overline{A_{i}}$, where $\overline{A_{i}}$ denotes the mean of the $i$-th column, and similar operations apply to other terms. With the centered distance matrix, the distance covariance is computed as $\mathrm{d}\operatorname{Cov}(x,z)=\frac{1}{\mathrm{n}^{2}}\sum_{\mathrm{i},\mathrm{j}}A_{ij}^{c}B_{ij}^{c}$, where $B$ represents the matrix of features $z$, and $n$ denotes the batch size.
@@
(13) $\mathcal{L}_{\text{Pearson }}=\frac{1}{C^{2}}\sum_{c_{1}=1}^{C}\sum_{c_{2}=1}^{C}\left(\operatorname{pCor}_{c_{1},c_{2}}\right)^{2}=\frac{1}{C^{2}}\sum_{c_{1}=1}^{C}\sum_{c_{2}=1}^{C}\left(\frac{\operatorname{Cov}_{c_{1},c_{2}}}{\sigma_{c_{1}}\cdot\sigma_{c_{2}}}\right)^{2}$
@@
Label smoothing. Finally, to ensure model usability, we draw inspiration from (Struppek et al., 2023) and apply LS to refine the target loss, as shown in formula (14). Additionally, KL divergence is used instead of cross-entropy (CE) as the primary loss function for optimization.
(14) $LS\left(\hat{y}_{c}\right)=1-\alpha+\frac{\alpha}{K},\quad LS\left(\hat{y}_{i}\right)=\frac{\alpha}{K}$
@@
CE enhances inter-class feature distinctions, while KL divergence supports smoother and more generalized feature extraction between classes (Struppek et al., 2023). Within-class feature extraction captures universal characteristics, ensuring generalization and robustness for the target task. This optimization reduces feature separability, increasing the difficulty of attack, as demonstrated in Section 6.3.
@@
6. Experiments
In this section, we systematically evaluate and compare the ability of SiftFunnel and existing approaches to resist MIA. The specific experimental setup and evaluation metrics are shown below.
6.1. Experimental Setup
1) Datasets. We used three types of image recognition datasets, with processing details outlined as follows. The detailed data allocation is shown in Table 1.
  • CIFAR-10 (Krizhevsky et al., 2009). The CIFAR-10 dataset consists of 60,000 RGB images categorized into ten classes with an original resolution of $32\times 32$. To make the results clearer and more intuitive, we used an unscaled resolution of $64\times 64$.
  • FaceScrub (Ng and Winkler, 2014). FaceScrub is a URL dataset with 100,000 images of 530 actors, which contains 265 male actors and 265 female actors. However, since not every URL was available during the writing period, we downloaded a total of 43,149 images for 530 individuals and resized the images to $64\times 64$.
  • CelebA (Liu et al., 2015). CelebA is a dataset with 202,599 images of 10,177 celebrities. We used the same crop as (Yang et al., 2019) to remove the background of images in this dataset other than faces to reduce the impact on the experiment. To eliminate individual overlap, we removed a total of 6,878 images of 296 individuals and similarly resized the images to $64\times 64$.
  • ChestX-ray (Sait et al., 2020). This dataset is a curated collection of COVID-19 chest X-ray images compiled from 15 publicly available datasets. It contains 1,281 COVID-19 X-rays, 3,270 normal X-rays, 1,656 viral pneumonia X-rays, and 3,001 bacterial pneumonia X-rays with a resolution of $128\times 128$.
Table 1. Data allocation of the target model and attack model.
@@ 6.1. Experimental Setup
Table 2. Quantitative evaluation of CNN on CIFAR-10 with different defenses against two MIAs. ↑ indicates higher is better, ↓ indicates lower is better. Bolded values are the best. Blue highlights show unprotected baseline performance, red indicates increased edge device load, yellow indicates reduced load, and gray shows gray-box defense performance.
MLE-based Attack (He et al., 2019) Gen-based Attack (He et al., 2020)
@@
Adding noise (He et al., 2020) ($\sigma$ = 0.8)
@@
NLS (Struppek et al., 2023) ($\alpha$ = -0.05)
87.60% (-0.68%)
0.0002 42.2684 0.9996 0.0014 33.3732 0.9949 0.0573 47,183 299,520
MID (Wang et al., 2021) (1e-2) 81.53% (-6.75%) 1.74e-5 50.4675 0.9999 0.0013 34.5184 0.9967 0.0584 45,198
BiDO (Peng et al., 2022) (2; 20) 84.12% (-4.16%) 0.6699 6.5111 0.0138 0.035 19.2965 0.8751 0.0429 3,345
VIB (Wang et al., 2024) 88.21% (-0.07%) 0.5552 7.3264 0.0358 0.0042 28.5054 0.9842 0.0551 18,069
IB-based InfoSCISSORS (Duan et al., 2023) 83.46% (-4.83%) 0.5715 7.2007 0.1345 0.0062 26.8799 0.9948 0.0549 43,364 299,520
@@
Figure 4. A visual evaluation of the attack effectiveness of two MIAs on CNN edge models protected by various defense techniques, with NND & IB-based results specifically reflecting the gray-box scenario.
[…]
4
@@
3) Target Edge Model and Implementation Details. We employ the same comprehensive model architecture as detailed in (Yang et al., 2019), comprising four CNN blocks succeeded by two fully connected layers. This model is trained using the Adam optimizer with a learning rate of 2e-4, a $\beta_{1}$ parameter of 0.5, and a ReduceLROnPlateaud scheduler with a reduction factor of 0.5 and patience of 25. Each CNN block is composed of a convolutional layer, a batch normalization layer, a max-pooling layer, and a ReLU activation function. The target edge model encompasses the initial two CNN blocks. For the training of the edge model protected by SiftFunnel, we have specified the parameters $\lambda_{1}$ to 3.5, $\alpha$ to 0.35, $\lambda_{2}$ to 0.8, and $\lambda_{3}$ to 0.6 in Equation (15). In addition, to argue the impact of the model architecture for the defense, we use different CNNs as follows: (1) VGG16 adapted from (Simonyan, 2014); (2) ResNet-18 adapted from (He et al., 2016), the learning rates are 3e-4. The division positions of ResNet and VGG16 are not fixed. For detailed accuracy and how to divide them, see Section 6.3 for analysis. All experiments were performed on two RTX 4090 GPUs and an Intel Core i9-14900KF ×$\times 32$ CPU.
@@
4) Comparison of Defense Methods. We aim to compare three existing categories of defense methods. For Perturbation-based defenses, we select Adding noise & Dropping features (He et al., 2020), DPSGD (Abadi et al., 2016), and negative label smoothing (NLS) (Struppek et al., 2023). For IB-based defenses, we choose MID (Wang et al., 2021), BiDO (with HSIC) (Peng et al., 2022), VIB (Wang et al., 2024), and InfoSCISSORS (Duan et al., 2023). For NND & IB-based defenses, we opt for AE-based protection (Azizian and Bajić, 2024). Our method falls into the NND & IB-based defense category, which also requires analyzing the defensive effect in a gray-box scenario. Furthermore, all comparison methods have been selected and set with appropriate hyper-parameters for comparison, with detailed settings found in Section 6.3. Lastly, all methods are trained without adversarial training, focusing solely on the effects of the original defense methods.
6.2. Evaluation Metrics
@@
We selected 7 evaluation metrics to assess the aspects of usability, privacy, and deployability. For usability assessment, we employ Test Accuracy (Test ACC), which evaluates the final performance of the model protected by defense techniques on the test dataset. Regarding privacy assessment, we initially utilize MSE, Peak Signal-to-Noise Ratio (PSNR), and Structural Similarity Index (SSIM) to evaluate the effectiveness of MIA on reconstructing the input data of the protected model (Yin et al., 2023). Here, a higher MSE and lower PSNR and SSIM (ranging from 0 to 1) indicate better defense effectiveness. In addition, we employ MINE (Belghazi et al., 2018), as mentioned in Section  4.3, to assess Mutual Information (MI) (ranging from 0 to 1) and the Effective Information Mean $\delta(z)$, which is the mean of the non-zero output by the edge model on the test dataset. The specific training methods and model architecture for MINE are referenced from https://github.com/SiftFunnel/SiftFunnel. For deployability, we consider the Edge model’s parameter count $|\theta_{\text{edge}}|$ and Inference Latency. The model parameter count directly determines memory or GPU memory occupation, while inference latency is tested on both CPU and GPU.
@@
Table 3. Quantitative evaluation of ResNet-18 on CIFAR-10 with different defenses against two MIAs. The target edge model is the first residual block of ResNet-18. ↑ indicates higher is better, ↓ indicates lower is better. Bolded values are the best. The meanings of color-coded regions are consistent with those in Table 2.
MLE-based Attack (He et al., 2019) Gen-based Attack (He et al., 2020)
@@
Adding noise (He et al., 2020) ($\sigma$ = 2)
@@
Table 4. Quantitative evaluation of different defense methods in protecting CNN against MIAs on the FaceScrub dataset. Changes in data types alter the inference task, requiring adjustments to the parameters of the defense methods.
MLE-based Attack (He et al., 2019) Gen-based Attack (He et al., 2020)
@@
NLS (Struppek et al., 2023) ($\alpha$ = -0.2)
87.97% (-0.34%) 0.0004 38.8904 0.9980 0.0018 32.3274 0.9906 0.0496 44,783 299,520
MID (Wang et al., 2021) (1e-2) 79.45% (-10.98%) 0.0017 32.3435 0.9897 0.0018 32.1285 0.9901 0.0507 39,277
BiDO (Peng et al., 2022) (2; 20) 88.02% (-0.29%) 0.0004 39.0838 0.9978 0.0018 32.2048 0.9904 0.0493 44,915
VIB (Wang et al., 2024) 88.05% (-0.26%) 0.7354 6.1060 0.0032 0.0033 29.5697 0.9823 0.0493 15,649
IB-based InfoSCISSORS (Duan et al., 2023) 81.61% (-6.70%) 0.7449 6.0504 0.0026 0.0060 26.9803 0.9675 0.0408 57,050 299,520
@@ 6.3. Experimental Results
1) The effect of various defenses on the MI and $\delta(z)$. As shown in Table 2, defense effectiveness is reflected by MI and $\delta(z)$, aligning with the criterion $D_{\text{mia}}$. Specifically, unprotected edge models are highly vulnerable to MIAs, with MI at 0.0566, MSE below 0.01, and SSIM exceeding 0.98. Perturbation-based defenses reduce MI by at most 11% but cause over a 10% drop in Test ACC, with NLS even increasing MI in shallow networks, weakening resistance to attacks. UIB-based defenses slightly reduce usability but effectively counter MLE-based MIAs, except for MID, which shows weaker resistance. Among them, BiDO significantly impacts MI and $\delta(z)$, providing effective defense, but its parameter selection and kernel width configuration (as shown in the table) limits generalizability, as confirmed in subsequent experiments. This experiment follows the InfoSCISSORS setup, ensuring CLUB receives a complete dataset and sufficient training for convergence. While CLUB effectively defends against MLE-based MIAs, it struggles against Gen-based MIAs and incurs a 4.83% accuracy loss. AE-based defenses improve model usability, likely due to enhanced model capacity, while MI reductions confirm reasonable defense effectiveness. However, gray-box scenarios weaken AE-based defenses, and their added complexity increases storage costs by 8×8$\times$8 ×, as seen in $|\theta_{\text{edge}}|$. SiftFunnel achieves the best defense performance against both MIAs while keeping accuracy loss below 3%. It achieves MSE of 0.6792 (MLE-based MIA) and above 0.1 (Gen-based MIA). In gray-box scenarios, it maintains MSE of 0.06 and SSIM of 0.7647, ensuring robust resistance. Furthermore, MI and $\delta(z)$ confirm its optimal impact on the key factors of $D_{\text{mia}}$. Additionally, $|\theta_{\text{edge}}|$ reveals a nearly 20×20$\times$20 × reduction in edge model load, highlighting SiftFunnel’s efficiency.
Table 5. SiftFunnel Ablation Study on CIFAR-10. The target edge model is the first residual block of ResNet-18, with Gen-based MIA as the attack method. The Red area indicates the strong attack hypothesis from Section 5.2.
@@
2) The effect of model architecture. Modifying the model architecture does not alter the parameter settings of the proposed method, which remain consistent with Section 6.1. We adapt the model to ResNet-18 with skip connections, using the first residual block as the edge model. This setup poses challenges for defense techniques, as shallow networks retain simple structures and do not modify spatial dimensions. Table 3 shows that unprotected models exhibit minimal MI impact, resulting in weak resistance to MIAs, with MLE-based MIA achieving an MSE of $10^{-4}$ and Gen-based MIA $10^{-3}$. Under these conditions, perturbation-based defenses fail to balance usability and privacy, leading to poor defense performance. Among IB-based defenses, only VIB demonstrates resistance to MLE-based MIA by optimizing the edge model through $I(x,z)$ estimation as a loss function, but its effectiveness remains limited, particularly against Gen-based MIA. For BiDO, despite applying kernel bandwidth estimation methods from (Peng et al., 2022) and extensive experimental tuning, it fails to achieve a configuration that effectively balances usability and privacy, resulting in suboptimal defense performance. InfoSCISSORS impacts MI under this configuration, but the increase in $\delta(z)$ undermines its effectiveness against Gen-based MIAs, leading to only marginal improvements. AE-based defenses show some resistance to attacks on low-dimensional features in earlier stages, but their effectiveness diminishes under the ResNet configuration. While they reduce $\delta(z)$ by 60%, they have minimal impact on MI, achieving only a 0.04% reduction. Consequently, AE-based defenses struggle to counter both black-box and gray-box Gen-based MIAs effectively.
Table 6. Quantitative evaluation of the resistance to MIA across different architectures on CIFAR-10. The edge model divisions for CNN and ResNet-18 follow the configurations in Tables 2 and 3, while for VGG16, the edge model is set at the first ReLU layer after a max-pooling operation.
MLE-based Attack (He et al., 2019) Gen-based Attack (He et al., 2020)
@@
In contrast, SiftFunnel reduces MI by 50% and lowers $\delta(z)$ to the millesimal scale, delivering the strongest resistance against MIAs while preserving usability and deployability. Notably, it achieves a minimum SSIM of 0.0793, underscoring its superior ability to balance privacy and usability. The different protection performances of the three model architectures with the same dataset are shown in Table 6.
@@
3) The effect of data type. Changes in data types lead to shifts in task requirements. For instance, in this study, the facial recognition task involves 530 classes, while ChestX-ray classification involves 4 classes, necessitating parameter adjustments as shown in Table 4. As analyzed in the previous section on model architectures, to achieve better defense performance, we adjusted the edge model divisions for ResNet-18 and VGG. Specifically, the division points were set after the second residual block for ResNet and after the first ReLU layer following the second max-pooling layer for VGG. The results, summarized in Table 4 and Figure 5, show that the defense performance remains unaffected by changes in data types. SiftFunnel continues to be the best defense method, offering an optimal balance between usability, privacy, and deployability.
@@
As shown in Figure 5, the VGG16 model protected by SiftFunnel demonstrates better resistance to Gen-based MIA in terms of the visual effects of usability and privacy on facial and X-ray data, with PSNR values consistently around 15.
@@
4) Ablation Study. To assess the effectiveness of SiftFunnel’s model structure, loss function design, and defense capability, we conducted ablation experiments using Gen-based MIA. As shown in Table 5, we separately removed the Information Funnel and Attention Blocks from the model structure. The results indicate that removing the Funnel had minimal impact on accuracy, reducing it by only 0.1%, whereas removing the attention mechanism significantly increased MI, weakening privacy protection. For the loss function, we examined the effect of removing KL divergence, linear and nonlinear constraints, and the $l_{1}$-norm. The results show that KL divergence and $l_{1}$-norm provided moderate defense improvements, while nonlinear constraints were critical—removing them increased MSE to 0.0268. Additionally, we validated the hypothesis from Section 5.2, demonstrating that even when adversaries reconstructed inputs using cloud-compensated features, SiftFunnel still outperformed existing defenses in protecting transmitted features.
@@
Figure 5. Quantitative and visual evaluation of privacy and usability against Gen-based MIA in the VGG16 edge model on the ChestX-ray ($\lambda_{1}=3.5,\lambda_{2}=1.2,\lambda_{3}=0.8$) and FaceScrub ($\lambda_{1}=4.5,\lambda_{2}=0.4,\lambda_{3}=0.1$).
[…]
5

### v5

@@ 5.3 The Design of Loss Function
Some defense approaches utilize Constrained Covariance (COCO) [52] to estimate correlation, focusing primarily on the linear relationship between $x$ and $z$. However, since neural networks predominantly perform nonlinear operations, studies have shown that this constraint is not only ineffective but may also lead to reduced usability [18]. Alternatively, some approaches employ the Hilbert-Schmidt Independency Criterion (HSIC) [54] to estimate nonlinear correlation. However, kernel-based calculations in HSIC rely heavily on the appropriate selection of $\sigma$. Improper $\sigma$ value can introduce bias in correlation estimation, limiting the generalizability of this method and reducing its effectiveness across various scenarios.
@@
$\mathcal{L}_{\text{d Cor}}=\mathrm{d}\operatorname{Cor}(x,z)=\frac{\mathrm{d}\operatorname{Cov}(x,z)}{\sqrt{\mathrm{d}\operatorname{Var}(x)\cdot\mathrm{d}\operatorname{Var}(z)}}$ (11)
@@
To compute the numerator, a distance matrix $A$ is first calculated, where each element is the Euclidean distance between samples, defined as $A=\left\|x_{i}-x_{j}\right\|_{2}$. To remove the influence of sample shifts, the distance matrix is centered as $A_{ij}^{c}=A_{ij}-\overline{A_{i}}-\overline{A_{j}}+\overline{A_{i}}$, where $\overline{A_{i}}$ denotes the mean of the $i$-th column, and similar operations apply to other terms. With the centered distance matrix, the distance covariance is computed as $\mathrm{d}\operatorname{Cov}(x,z)=\frac{1}{\mathrm{n}^{2}}\sum_{\mathrm{i},\mathrm{j}}A_{ij}^{c}B_{ij}^{c}$, where $B$ represents the matrix of features $z$, and $n$ is the batch size.
@@
$\mathcal{L}_{\text{Pearson }}$ $=\frac{1}{C^{2}}\sum_{c_{1}=1}^{C}\sum_{c_{2}=1}^{C}\left(\operatorname{pCor}_{c_{1},c_{2}}\right)^{2}$ (12)
$=\frac{1}{C^{2}}\sum_{c_{1}=1}^{C}\sum_{c_{2}=1}^{C}\left(\frac{\operatorname{Cov}_{c_{1},c_{2}}}{\sigma_{c_{1}}\cdot\sigma_{c_{2}}}\right)^{2}$
@@
Label smoothing. Finally, to ensure model usability, we draw inspiration from [33] and apply LS to refine the target loss, as shown in formula (14). Additionally, KL divergence is used instead of cross-entropy (CE) as the primary loss function for optimization.
$LS\left(\hat{y}_{c}\right)=1-\alpha+\frac{\alpha}{K},\quad LS\left(\hat{y}_{i}\right)=\frac{\alpha}{K}$ (13)
@@
While the $z$ at this layer have not fully transitioned into decision-level semantics, the combination of architectural constraints and the implicit regularization of $I(x,z)$ imposed by the loss function encourages more task-relevant representations. Consequently, positive-factor LS can induce slight intra-class compactness, modestly reducing the feature entropy and hindering inversion attacks [33], as evidenced in Section 6.3.
Even in cases where LS increases entropy due to insufficient decision-oriented expression, its impact on attack vulnerability remains limited. Ultimately, label smoothing is intended to promote generalization, with improved resistance to MIAs appearing only as a byproduct when sufficient constraints are imposed on the feature representations.
@@
6 Experiments
6.1 Experimental Setup
1) Datasets. We used three types of image recognition datasets, with processing details outlined as follows. The detailed data allocation is shown in Table I.
  • CIFAR-10 [46]. The CIFAR-10 dataset consists of 60,000 RGB images categorized into ten classes with an original resolution of $32\times 32$. To make the results clearer and more intuitive, we used an unscaled resolution of $64\times 64$.
  • FaceScrub [55]. FaceScrub is a URL dataset with 100,000 images of 530 actors, which contains 265 male actors and 265 female actors. However, since not every URL was available during the writing period, we downloaded a total of 43,149 images for 530 individuals and resized the images to $64\times 64$.
  • CelebA [56]. CelebA is a dataset with 202,599 images of 10,177 celebrities. We used the same crop as [10] to remove the background of images in this dataset other than faces to reduce the impact on the experiment. To eliminate individual overlap, we removed a total of 6,878 images of 296 individuals and similarly resized the images to $64\times 64$.
  • ChestX-ray [57]. This dataset is a curated collection of COVID-19 chest X-ray images compiled from 15 publicly available datasets. It contains 1,281 COVID-19 X-rays, 3,270 normal X-rays, 1,656 viral pneumonia X-rays, and 3,001 bacterial pneumonia X-rays with a resolution of $128\times 128$.
TABLE I: Data allocation of the target model and attack model.
@@ 6.1 Experimental Setup
TABLE II: Quantitative evaluation of CNN on CIFAR-10 with different defenses against two MIAs. ↑ indicates higher is better, ↓ indicates lower is better. Bolded values are the best. Blue highlights show unprotected baseline performance, red indicates increased edge device load, yellow indicates reduced load, and gray shows gray-box defense performance.
MLE-based Attack [6] Gen-based Attack [16]
@@
Adding noise [16] ($\sigma$ = 0.8)
@@
NLS [33] ($\alpha$ = -0.05)
85.53% (-2.75%)
0.0002 42.2684 0.9996 0.0015 33.0003 0.9943 0.0573 47,183 299,520
MID [21] (1e-2) 81.53% (-6.75%) 1.74e-5 50.4675 0.9999 0.0013 34.5184 0.9967 0.0584 45,198
BiDO [18] (2; 20) 84.12% (-4.16%) 0.6699 6.5111 0.0138 0.035 19.2965 0.8751 0.0429 3,345
VIB [19] 88.21% (-0.07%) 0.5552 7.3264 0.0358 0.0042 28.5054 0.9842 0.0551 18,069
IB-based InfoSCISSORS [34] 83.46% (-4.83%) 0.5715 7.2007 0.1345 0.0062 26.8799 0.9948 0.0549 43,364 299,520
@@
Figure 4: A visual evaluation of the attack effectiveness of two MIAs on CNN edge models protected by various defense techniques, with NND & IB-based results specifically reflecting the gray-box scenario.
@@
3) Target Edge Model and Implementation Details. We employ the same comprehensive model architecture as detailed in [10], comprising four CNN blocks succeeded by two fully connected layers. This model is trained using the Adam optimizer with a learning rate of 2e-4, a $\beta_{1}$ parameter of 0.5, and a ReduceLROnPlateaud scheduler with a reduction factor of 0.5 and patience of 25. Each CNN block is composed of a convolutional layer, a batch normalization layer, a max-pooling layer, and a ReLU activation function. The target edge model encompasses the initial two CNN blocks. For the training of the edge model protected by SiftFunnel, we have specified the parameters $\lambda_{1}$ to 3.5, $\alpha$ to 0.35, $\lambda_{2}$ to 0.8, and $\lambda_{3}$ to 0.6 in Equation (15). In addition, to argue the impact of the model architecture for the defense, we use different CNNs as follows: (1) VGG16 adapted from [58]; (2) ResNet-18 adapted from [38], the learning rates are 3e-4. The division positions of ResNet and VGG16 are not fixed. For detailed accuracy and how to divide them, see Section 6.3 for analysis. All experiments were performed on two RTX 4090 GPUs and an Intel Core i9-14900KF ×$\times 32$ CPU.
@@
4) Comparison of Defense Methods. We aim to compare three existing categories of defense methods. For Perturbation-based defenses, we select Adding noise & Dropping features [16], DPSGD [59], and negative label smoothing (NLS) [33]. For IB-based defenses, we choose MID [21], BiDO (with HSIC) [18], VIB [19], and InfoSCISSORS [34]. For NND & IB-based defenses, we opt for AE-based protection [15]. Our method falls into the NND & IB-based defense category, which also requires analyzing the defensive effect in a gray-box scenario. Furthermore, all comparison methods have been selected and set with appropriate hyper-parameters for comparison, with detailed settings found in Section 6.3. Lastly, all methods are trained without adversarial training, focusing solely on the effects of the original defense methods.
6.2 Evaluation Metrics
@@
We selected 7 evaluation metrics to assess the aspects of usability, privacy, and deployability. For usability assessment, we employ Test Accuracy (Test ACC), which evaluates the final performance of the model protected by defense techniques on the test dataset. Regarding privacy assessment, we initially utilize Mean Squared Error (MSE), Peak Signal-to-Noise Ratio (PSNR), and Structural Similarity Index (SSIM) to evaluate the effectiveness of MIA on reconstructing the input data of the protected model [8]. Here, a higher MSE and lower PSNR and SSIM (ranging from 0 to 1) indicate better defense effectiveness. In addition, we employ MINE [24], as mentioned in Section  4.3, to assess Mutual Information (MI) (ranging from 0 to 1) and the Effective Information Mean $\delta(z)$, which is the mean of the non-zero output by the edge model on the test dataset. The specific training methods and model architecture for MINE are referenced from https://github.com/SiftFunnel/SiftFunnel. For deployability, we consider the Edge model’s parameter count $|\theta_{\text{edge}}|$ and Inference Latency. Model parameters directly affect memory usage, and inference latency is measured on both CPU and GPU.
@@
TABLE III: Quantitative evaluation of ResNet-18 on CIFAR-10 with different defenses against two MIAs. The target edge model is the first residual block of ResNet-18. ↑ indicates higher is better, ↓ indicates lower is better. Bolded values are the best. The meanings of color-coded regions are consistent with those in Table II.
MLE-based Attack [6] Gen-based Attack [16]
@@
Adding noise [16] ($\sigma$ = 2)
@@
TABLE IV: Quantitative evaluation of different defense methods in protecting CNN against MIAs on the FaceScrub dataset. Changes in data types alter the inference task, requiring adjustments to the parameters of the defense methods.
MLE-based Attack [6] Gen-based Attack [16]
@@
NLS [33] ($\alpha$ = -0.2)
88.03% (-0.28%) 0.0004 38.8904 0.9980 0.0017 32.5480 0.9911 0.0515 44,783 299,520
MID [21] (1e-2) 79.45% (-10.98%) 0.0017 32.3435 0.9897 0.0018 32.1285 0.9901 0.0507 39,277
BiDO [18] (2; 20) 88.02% (-0.29%) 0.0004 39.0838 0.9978 0.0018 32.2048 0.9904 0.0493 44,915
VIB [19] 88.05% (-0.26%) 0.7354 6.1060 0.0032 0.0033 29.5697 0.9823 0.0493 15,649
IB-based InfoSCISSORS [34] 81.61% (-6.70%) 0.7449 6.0504 0.0026 0.0060 26.9803 0.9675 0.0408 57,050 299,520
@@ 6.3 Experimental Results
1) The effect of various defenses on the MI and $\delta(z)$. As shown in Table II, defense effectiveness is reflected by MI and $\delta(z)$, aligning with the criterion $D_{\text{mia}}$. Specifically, unprotected edge models are highly vulnerable to MIAs, with MI at 0.0566, MSE below 0.01, and SSIM exceeding 0.98. Perturbation-based defenses reduce MI by at most 11% but cause over a 10% drop in Test ACC, with NLS even increasing MI in shallow networks, weakening resistance to attacks. From the theoretical analysis, it can be seen that DPSGD does not alter the complexity of the mapping from $x$ to $z$, nor does it directly affect the intermediate feature $z$ itself. As a result, it reduces model accuracy without effectively mitigating MIA risks in the CI setting. IB-based defenses slightly reduce usability but effectively counter MLE-based MIAs, except for MID, which shows weaker resistance. Among them, BiDO significantly impacts MI and $\delta(z)$, providing effective defense, but its parameter selection and kernel width configuration (as shown in the table) limits generalizability, as confirmed in subsequent experiments. This experiment follows the InfoSCISSORS setup, ensuring CLUB receives a complete dataset and sufficient training for convergence. While CLUB effectively defends against MLE-based MIAs, it struggles against Gen-based MIAs and incurs a 4.83% accuracy loss. AE-based defenses improve model usability, due to enhanced model capacity, while MI reductions confirm reasonable defense effectiveness. However, gray-box scenarios weaken AE-based defenses, and their added complexity increases storage costs by 8×8$\times$8 ×, as seen in $|\theta_{\text{edge}}|$. SiftFunnel achieves the best defense performance against both MIAs while keeping accuracy loss below 3%. It achieves MSE of 0.6792 (MLE-based MIA) and above 0.1 (Gen-based MIA). In gray-box scenarios, it maintains MSE of 0.06 and SSIM of 0.7647, ensuring robust resistance. Furthermore, MI and $\delta(z)$ confirm its optimal impact on the key factors of $D_{\text{mia}}$. Additionally, $|\theta_{\text{edge}}|$ reveals a nearly 20×20$\times$20 × reduction in edge model load, highlighting SiftFunnel’s efficiency.
TABLE V: SiftFunnel Ablation Study on CIFAR-10. The target edge model is the first residual block of ResNet-18, with Gen-based MIA as the attack method. The Red area indicates the strong attack hypothesis from Section 5.2.
@@
2) The effect of model architecture. Modifying the model architecture does not alter the parameter settings of the proposed method, which remain consistent with Section 6.1. We adapt the model to ResNet-18 with skip connections, using the first residual block as the edge model. This setup poses challenges for defense techniques, as shallow networks retain simple structures and do not modify spatial dimensions. Table III shows that unprotected models exhibit minimal MI impact, resulting in weak resistance to MIAs, with MLE-based MIA achieving an MSE of $10^{-4}$ and Gen-based MIA $10^{-3}$. Under these conditions, perturbation-based defenses fail to balance usability and privacy, leading to poor defense performance. Among IB-based defenses, only VIB demonstrates resistance to MLE-based MIA by optimizing the edge model through $I(x,z)$ estimation as a loss function, but its effectiveness remains limited, particularly against Gen-based MIA. For BiDO, despite applying kernel bandwidth estimation methods from [18] and extensive experimental tuning, it fails to achieve a configuration that effectively balances usability and privacy, resulting in suboptimal defense performance. InfoSCISSORS impacts MI under this configuration, but the increase in $\delta(z)$ undermines its effectiveness against Gen-based MIAs, leading to only marginal improvements. AE-based defenses show some resistance to attacks on low-dimensional features in earlier stages, but their effectiveness diminishes under the ResNet configuration. While they reduce $\delta(z)$ by 60%, they have minimal impact on MI, achieving only a 0.04% reduction. Consequently, AE-based defenses struggle to counter both black-box and gray-box Gen-based MIAs effectively.
TABLE VI: Quantitative evaluation of MIA resistance on CIFAR-10 across different architectures. Edge models for CNN and ResNet-18 follow Tables II and III; for VGG16, the split is placed at the first ReLU after max pooling.
MLE-based Attack [6] Gen-based Attack [16]
@@
In contrast, SiftFunnel reduces MI by 50% and lowers $\delta(z)$ to the millesimal scale, offering the strongest MIA resistance while maintaining usability. It achieves a minimum SSIM of 0.0793, highlighting its privacy–utility balance. Table VI compares performance across architectures.
@@
3) The effect of data type. Changes in data types lead to shifts in task requirements. For instance, in this study, the facial recognition task involves 530 classes, while ChestX-ray classification involves 4 classes, necessitating parameter adjustments as shown in Table IV. As analyzed in the previous section on model architectures, to achieve better defense performance, we adjusted the edge model divisions for ResNet-18 and VGG. Specifically, the division points were set after the second residual block for ResNet and after the first ReLU layer following the second max-pooling layer for VGG. The results, summarized in Table IV and Figure 5, show that the defense performance remains unaffected by changes in data types. SiftFunnel continues to be the best defense method, offering an optimal balance between usability, privacy, and deployability.
@@
As shown in Figure 5, the VGG16 model with SiftFunnel shows better resistance to Gen-based MIA in terms of the visual effects of usability and privacy on facial and X-ray data, with PSNR values consistently around 15.
@@
4) Ablation Study. To assess the effectiveness of SiftFunnel’s model structure, loss function design, and defense capability, we conducted ablation experiments using Gen-based MIA. As shown in Table V, we separately removed the Information Funnel and Attention Blocks from the model structure. The results indicate that removing the Funnel had minimal impact on accuracy, reducing it by only 0.1%, whereas removing the attention mechanism significantly increased MI, weakening privacy protection. For the loss function, we examined the effect of removing KL divergence, linear and nonlinear constraints, and the $l_{1}$-norm. The results show that KL divergence and $l_{1}$-norm provided moderate defense improvements, while nonlinear constraints were critical—removing them increased MSE to 0.0268. Additionally, we validated the hypothesis from Section 5.2, demonstrating that even when adversaries reconstructed inputs using cloud-compensated features, SiftFunnel still outperformed existing defenses in protecting transmitted features.
@@
Figure 5: Quantitative and visual evaluation of privacy and usability against Gen-based MIA in the VGG16 edge model on the ChestX-ray ($\lambda_{1}=3.5,\lambda_{2}=1.2,\lambda_{3}=0.8$) and FaceScrub ($\lambda_{1}=4.5,\lambda_{2}=0.4,\lambda_{3}=0.1$).